# Combining Crown Structures for Vulnerability Measures

### v1

@@ 4 Kernels for Component Order Connectivity
In this section, we consider COC, a refined version of VI. Given a graph $G=(V,E)$ and parameters $k,W\in\mathds{N}$, COC aims to remove at most $k$ vertices so that resulting components have at most $W$ vertices each. We present a kernelization algorithm that provides a $2kW$ vertex-kernel in an FPT-runtime when parameterized by the size of a maximum ($W+1$)-packing, as stated in the following theorem.
@@
A kernel of size $2kW$ is also presented in [DBLP:conf/iwpec/KumarL16] with an exponential runtime using linear programming methods, which however, as mentioned above, can be improved to an FPT-runtime in the parameter $W$. In contrast, our result is fully combinatorial and has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$. Moreover, we showcase how our algorithm transforms into a polynomial one for two cases: when $W=1$ (Vertex Cover) and for claw-free graphs. While generalizing our FPT-runtime algorithm into polynomial time eludes us, our understanding of crown decompositions structural properties holds promise for future progress.
Before, i.e. for VI, wVI, and wCOC, crown structures were integrated into the head and crown of the BCD. In contrast, we now incorporate crown structures into the body of the BCD. Of particular interest are the so-called strictly reducible pairs introduced by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] who prove that such a structure must exist in graphs with more than $2kW$ vertices.
@@
We say $(A,B)$ is a minimal (strictly) reducible pair if there does not exist a (strictly) reducible pair $(A^{\prime},B^{\prime})$ with $A^{\prime}\subset A$ and $B^{\prime}\subseteq B$. A (strictly) reducible pair is basically a weighted crown decomposition where $A$ is the head and $B$ is the crown.
@@
In essence, the kernelization algorithm outlined below focuses on analyzing pairs $(A,B)$ in a $W$-BCD. It turns out that the head vertices $A$ have specific traits in a $W$-BCD, making it possible to locate them. Structurally, the algorithm operates as a bounded search tree, hence it is presented recursively. It takes an instance $(G,k,W)$ of COC as input, where $|V(G)|>2kW$ and each component of $G$ has at least $W+1$ vertices. The size limits for the input graph are not arbitrary; if $|V(G)|\leq 2kW$, there is nothing to do, and removing components smaller than $W+1$ is a safe reduction.
@@
For $V^{\prime}\subseteq V$ we define $\texttt{sep}_{W}(V^{\prime})\in\mathds{N}$ as the cardinality of a minimum $W$-separator in $G[V^{\prime}]$, and $\texttt{arg\hskip 2.84544ptsep}_{W}(V^{\prime})\subset V$ as an argument suitable to this cardinality. For a graph $G^{\prime}\subseteq G$ let $G^{>W}(G^{\prime})$ be the graph obtained by removing all components of size at most $W$ from $G$. Lastly, for a $W$-BCD $(C,H,\mathfrak{R},f)$ we define :=$\mathfrak{R}^{\prime}:=\{R\in\mathfrak{R}\mid|R|>2W\text{ and }\texttt{sep}_{W}(R)=1\}$ and $S_{\mathfrak{R}^{\prime}}:=\bigcup_{R\in\mathfrak{R}^{\prime}}\texttt{arg\hskip 2.84544ptsep}_{W}(R)$. (We show in LABEL:lemma::UniqueSep the uniqueness of $\texttt{arg\hskip 2.84544ptsep}_{W}(R)$ for $R\in\mathfrak{R}^{\prime}$).
@@
    Compute a a $W$-BCD $(C,H,\mathfrak{R},f)$ in $G$ and initialize $t=|H|+|\mathfrak{R}|$ and $S=\varnothing$.
@@
      Let $\mathcal{Q}$ be the connected components of size at most $W$ in $G-S$. Let $A=S\cup H$ and $B=C\cup V(\mathcal{Q})$. Compute a DBE $(A_{1},A_{2},W,f,2W-1)$ in $G[A\cup B]$ by using Theorem 2.2. If $A_{1}\neq\varnothing$, then terminate the algorithm and return $(G-(A_{1}\cup f^{-1}(A_{1})),k-|A_{1}|,W)$.
@@
    If the depth of the recursion is more than $\min(3t,k)$, then return.
@@
The interesting part of the algorithm is the localization of a minimal strictly reducible pair if it exists in the graph. Let $(A,B)$ be a minimal strictly reducible pair in $G$. As already mentioned, algorithm
@@
2kw is basically a bounded search tree with a working vertex set $S$, which are potential head vertices. The crucial step is ensuring that $S$ consistently matches the vertex set $A$. Once this alignment is achieved, localization of $(A,B)$ is assured, as step 2b (specifically Theorem 2.2) guarantees its extraction.
We conclude this part of the section by presenting a crucial lemma along with its implication for a depth-wise progression within the bounded search tree towards $S=A$ in algorithm

### v2

@@ 4 Kernels for Component Order Connectivity
In this section, we consider COC, a refined version of VI. For VI, wVI, and wCOC, crown structures were integrated into the head and crown of the BCD. In contrast, we now incorporate them into the body of the BCD.
Given a graph $G=(V,E)$ and parameters $k,W\in\mathbb{N}$, COC is the problem of removing at most $k$ vertices so that each resulting component has at most $W$ vertices. We present a kernelization algorithm (
Algorithm 1.
2kw
) that produces a $2kW$-vertex kernel in FPT-runtime, parameterized by the size of a maximum $(W+1)$-packing, as stated in the following theorem.
@@
A kernel of size $2kW$ is also presented in [DBLP:conf/iwpec/KumarL16], however with an FPT-runtime in the parameter $W$ and using linear programming methods. In contrast, our result has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$ and is fully combinatorial. Although generalizing our FPT-runtime algorithm to polynomial time remains challenging, our insights into the structural properties of crown decompositions suggest promising directions for future work. Moreover, we demonstrate how our algorithm (AlgCOC-2) becomes polynomial for two cases: when $W=1$ (Vertex Cover) and for claw-free graphs, as stated in the theorems below, whose proofs can be found in Section LABEL:sec:kernels_for_special_COC.
Theorem 4.38.
For $W=1$, i.e. for the vertex cover problem, algorithm AlgCOC-2 provides a $2k$ vertex kernel in polynomial time.
Theorem 4.39.
The component order connectivity problem admits a $2kW$ vertex-kernel on claw-free graphs in polynomial time.
We begin with preliminaries on the reducible structures of COC and their properties, which are essential for our proofs. Following this, we focus on the kernelization algorithms and provide proofs of its correctness.
4.1 Reducible Structures of the Component Order Connectivity Problem and its Properties
We call a (not necessarily minimal) set of vertices that separates a graph into components of size at most $W$ a $W$-separator. We define $(G,k,W)$ as an instance of COC, where it is labeled as a yes-instance if there exists a $W$-separator of size $k$; otherwise, it is labeled as a no-instance. To design the kernelization algorithm, we briefly review, in addition to the preliminary content in Section 2, some definitions and theorems from [DBLP:conf/gecco/Baguley0N0PZ23, DBLP:conf/esa/Casel0INZ21, fomin2019kernelization, DBLP:conf/iwpec/KumarL16]. Of particular interest are the strictly reducible pairs introduced by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16], who proved that such structures must exist in graphs with more than $2kW$ vertices.
@@ 4.1 Reducible Structures of the Component Order Connectivity Problem and its Properties
We say that $(A,B)$ is a minimal (strictly) reducible pair if there does not exist a (strictly) reducible pair $(A^{\prime},B^{\prime})$ with $A^{\prime}\subset A$ and $B^{\prime}\subseteq B$. Clearly, if a graph $G$ contains a (strictly) reducible pair, then it also contains a minimal one. Furthermore, for $A^{\prime}\subseteq A$, we define $\mathbb{CC}(G[B])_{A^{\prime}}\subseteq\mathbb{CC}(G[B])$ as the set of components $\{Q\in\mathbb{CC}(G[B])\mid N(Q)\cap A^{\prime}\neq\varnothing\}$. Before discussing the connection between a (strictly) reducible pair and COC, we state a property of minimal strictly reducible pairs found by Baguley et al. [DBLP:conf/gecco/Baguley0N0PZ23].
Lemma 4.41 ([DBLP:conf/gecco/Baguley0N0PZ23], Lemma 4.16).
Let $(A,B)$ be a minimal strictly reducible pair in $G$ with parameter $W$. Then, for every $A^{\prime}\subseteq A$ we have $|V(\mathbb{CC}(G[B])_{A^{\prime}})|>|A^{\prime}|(2W-1)$.
Considering Lemma 4.41, if there is a minimal strictly reducible pair $(A,B)$ in $G$ and we remove a proper subset $A^{\prime}$ of $A$ from $G$, then $G-A^{\prime}$ still contains a strictly reducible pair with head vertices $A\setminus A^{\prime}$. Specifically, the specified vertex, guaranteed to receive a fractional assignment of at least $2W$ over $g$ from $\mathbb{CC}(G[B])$, can be chosen arbitrarily (cf. Definition 4.40). In our algorithm, these properties enable a stepwise approach in localizing minimal strictly reducible pairs.
A reducible pair $(A,B)$ is, in fact, a $(W,W)$-CD, or equivalently, an $(A,\varnothing,W,f,2W-1)$ DBE where $B=f^{-1}(A)$. Note that $|B|>|A|(2W-1)$; hence, it can be identified in graphs resembling bipartite graphs with the required expansion properties (see Definition 1 and Theorem 1). Here, the head and crown correspond to $A$ and $B$, respectively.
Once a $(W,W)$-CD is identified, we can reduce the COC instance by including the head vertices in the solution. In this case, no additional vertices from the crown are required, since after removing the head vertices, the crown’s connected components are isolated from $G$ and each connected component has size at most $W$.
In the following, we outline why this reduction is safe for COC. Let $P_{1},\dots,P_{m}\subseteq V$ be a $(W+1)$-packing. For a $W$-separator $S\subseteq V$, it holds that $S\cap P_{i}\neq\varnothing$ for all $i\in[m]$. Therefore, the size of a $(W+1)$-packing provides a lower bound on the number of vertices required for a $W$-separator. Considering a $(W,W)$-CD $C,H,f$ in $G$, the function $f$ enables the construction of a $(W+1)$-packing of size $|H|$ in $G[H\cup C]$. This essentially provides a lower bound of $|H|$ vertices for a $W$-separator in $G[H\cup C]$. On the other hand, $H$ serves as a $W$-separator of $G[A\cup B]$ while separating $C$ from the rest of the graph. These properties support the following theorem for $(W,W)$-CDs.
Lemma 4.42 ([DBLP:conf/esa/Casel0INZ21] Lemma 49).
Let $(G,k)$ be an instance of COC and $(A,B,f)$ a $(W,W)$-CD in $G$. Then, $(G,k)$ is a yes-instance if and only if $(G-(A\cup B),k-|A|)$ is a yes-instance.
Finally, the relationship between the existence of a strictly reducible pair and the attainment of the desired kernelization of $2kW$ is demonstrated in the following lemma.
Lemma 4.43 ([fomin2019kernelization], Lemma 6.14).
Let $(G,k,W)$ be an instance of the $W$-separator problem, such that each component in $G$ has size at least $W+1$. If $|V|>2kW$ and $(G,k,W)$ is a yes-instance, then there exists a strictly reducible pair $(A,B)$ in $G$.
The fundamental structure utilized in the kernelization algorithm is again the BCD. The established relationships to COC are presented in the following lemma.
Lemma 4.44 ([DBLP:conf/esa/Casel0INZ21], Lemma 48 and 49).
Let $(G,k,W)$ be an instance of COC, and $(C,H,\mathfrak{R},f)$ a $W$-BCD in $G$. $(G,k)$ is a yes-instance if and only if $(G-(H\cup C),k-|A|)$ is a yes-instance. Furthermore, if $|\mathfrak{R}|+|H|>k$, then $(G,k)$ is a no-instance.
4.2 Kernelization Algorithm
@@
In essence, the kernelization algorithm focuses on analyzing pairs $(A,B)$ in a $W$-BCD. We show that the head vertices $A$ have specific traits in a $W$-BCD, making it possible to locate them. Structurally, the algorithm operates as a bounded search tree, hence it is presented recursively. The input is an instance $(G,k,W)$ of COC, where $|V(G)|>2kW$ and each component of $G$ has at least $W+1$ vertices. The size limits for the input graph are not arbitrary; if $|V(G)|\leq 2kW$, there is nothing to do, and removing components smaller than $W+1$ is a safe reduction.
@@
For $V^{\prime}\subseteq V$ we define $\texttt{sep}_{W}(V^{\prime})\in\mathds{N}$ as the cardinality of a minimum $W$-separator in $G[V^{\prime}]$, and $\texttt{arg\hskip 2.84544ptsep}_{W}(V^{\prime})\subset V$ as an argument suitable to this cardinality. For a graph $G^{\prime}\subseteq G$ let $G^{>W}(G^{\prime})$ be the graph obtained by removing all components of size at most $W$ from $G$. Lastly, for a $W$-BCD $(C,H,\mathfrak{R},f)$ we define :=$\mathfrak{R}^{\prime}:=\{R\in\mathfrak{R}\mid|R|>2W\text{ and }\texttt{sep}_{W}(R)=1\}$ and $S_{\mathfrak{R}^{\prime}}:=\bigcup_{R\in\mathfrak{R}^{\prime}}\texttt{arg\hskip 2.84544ptsep}_{W}(R)$ (uniqueness of $\texttt{arg\hskip 2.84544ptsep}_{W}(R)$ for $R\in\mathfrak{R}^{\prime}$ is shown in LABEL:lemma::UniqueSep).
@@
    Compute a $W$-BCD $(C,H,\mathfrak{R},f)$ in $G$ and initialize $t=|H|+|\mathfrak{R}|$ and $S=\varnothing$.
@@
      Let $\mathcal{Q}$ be the connected components of size at most $W$ in $G-S$. Let $A=S\cup H$ and $B=C\cup V(\mathcal{Q})$. Compute a DBE $(A_{1},A_{2},W,f,2W-1)$ in $G[A\cup B]$ using Theorem 1. If $A_{1}\neq\varnothing$, then terminate the algorithm and return $(G-(A_{1}\cup f^{-1}(A_{1})),k-|A_{1}|,W)$.
@@
    If the depth of the recursion is more than $\min(3t,k)$, then break.
@@
We split the correctness proof of Algorithm
@@
2kw into the following three lemmas:
Lemma 4.45.
Let $(G=(V,E),k,W)$ be a yes-instance and let $(A,B)$ be a minimal strictly reducible pair in $G$. Then, Algorithm
Lemma 4.46.
In the Algorithm
Lemma 4.47.
If Algorithm
If Lemma 4.45 holds, then the instance can be reduced by algorithm